# Safety-guaranteed and Goal-oriented Semantic Sensing, Communication, and Control for Robotics

### 13 Mar 2026 v1

@@ VI Conclusion
We proposed a safety-guaranteed and goal-oriented semantic communication solution to maximize the robotic task effectiveness while guaranteeing safety requirements in wirelessly-connected robotic system. We first summarized the safety requirements and effectiveness metrics across representative robotic tasks, including robot arm grasping, unmanned aerial vehicle (UAV)-assisted tasks, and multi-robot exploration. We then systematically analyzed key safety and effectiveness challenges in sensing, communication, and control. Based on these, we further presented potential goal-oriented semantic-based solutions. Our UAV target tracking case study validates that our proposed solutions can significantly improve safety rate and tracking success rate by more than 2 times and 4.5 times, respectively.

### 26 Apr 2026 v2

@@ VI Conclusion
We investigated how to enable safety-aware goal-oriented semantic (SA-GS) sensing, communication, and control co-design in wirelessly-connected robotic systems, with the objective of maximizing task effectiveness subject to practical safety requirements. We first introduced the architecture and representative use cases of wirelessly-connected robotic systems. We then summarized the general safety requirements and effectiveness metrics across these use cases. Next, we systematically analyzed the key safety and effectiveness challenges in sensing, communication, and control. Based on this analysis, we presented potential SA-GS directions. Finally, a UAV target-tracking case study demonstrated that these directions can improve the safety rate and tracking success rate by more than 2 times and 4.5 times, respectively.